evaluate without using imaginary numbers:
fourth root of minus 36
fourth root of -36

Answer by Alan3354(69443)

evaluate without using imaginary numbers:
fourth root of minus 36
-----------
z^4 = -36
z^4 = 36cis(180)
z = sqrt(6)cis(45)
z = sqrt(6)cis(135)
z = sqrt(6)cis(225)
z = sqrt(6)cis(315)
---------